OpenStudy (anonymous):

The area of a rectangle is 102 square inches and the perimeter of the rectangle is 46 inches. Find the length and the width.

OpenStudy (anonymous):

A=LW P=2L+2W

OpenStudy (anonymous):

A=l*h per=2l+2h ---- so 102=l*h, 46= 2l+2h solve that ^

OpenStudy (anonymous):

{h = 17, l = 6} or {h = 6, l = 17}